# SRIF: Semantic Shape Registration Empowered by Diffusion-based Image Morphing and Flow Estimation

### v1

@@ Abstract.
* Equal contribution, $\dagger$ Corresponding to Ruqi Huang (ruqihuang@sz.tsinghua.edu.cn)
submissionid: 199ccs: Computing methodologies Shape analysis
@@ 1. Introduction
Estimating dense correspondences between 3D shapes serves as a cornerstone in many applications of computer graphics, including 3D reconstruction [338], animation [280] and statistical shape analysis [13], to name a few. Regarding shapes undergoing rigid or isometric deformations, the prior shape registration/matching techniques [226; 12; 41; 212] have laid down solid foundations on both theoretical and practical fronts. In this paper, we consider the problem of estimating semantically meaningful dense correspondences between shapes undergoing more general and complicated deformations.
@@
In the absence of a compact deformation prior, the purely geometric methods typically take a coarse-to-fine approach. Namely, one leverages geometric features to locate a small set of landmarks on both shapes, estimates sparse landmark correspondences, and finally propagates dense correspondences via minimizing distortions such as conformal [144], elastic energy [82]. It is worth noting, though, that the sparse correspondences derived from geometry are not necessarily relevant to semantics. As shown in qualitative results in Sec. 4, the resulting maps can suffer from such discrepancy, especially in the presence of significant heterogeneity.
@@
To this end, another line of works concentrate on producing high-quality semantic correspondences at the cost of dependency on user-defined landmarks [10; 184; 90; 263; 328]. Treating landmarks as anchor points, they cast the problem as correspondence interpolation, which can be conducted by optimizing certain geometric distortions. However, the dependency on manual annotations limits the practical utility of these works. Apart from hindering automation, insufficient annotation strength can lead to sub-optimal results as shown in Fig. 2.
@@
Figure 2. Maps obtained by state-of-the-art compatible remeshing techniques [328; 263] based on varying number of landmark correspondences. The absence of landmarks at local region can lead to erroneous matching (see the red boxes). On the other hand, our method delivers high-quality maps with no landmark annotation.
@@
More recently, the emergence of learning-based techniques has enabled data-driven semantic information extraction. In essence, one can learn canonical structural information from a collection of 3D shapes, which can help to match challenging non-rigid shapes [282; 84], or heterogeneous man-made objects [71; 344]. Unfortunately, limited by the amount of available 3D data, the above approaches are typically category-specific, weakening their practical utility.
@@
On the other hand, boosted by large-scale natural image datasets such as LAION [264], large vision (or vision-language) models (LVMs)  [210; 254; 239; 209] has attracted considerable attention from the community of 3D shape analysis  [3; 2; 202; 318]. The shared protocol is to project 3D shapes into multi-view 2D images. The latter can then be fed into LVMs for semantic encoding, which is finally aggregated back to 3D shapes for the respective task. While the above approach has been proven simple yet effective, we observe that 1) the distilled semantic information is in general coarse (e.g., segmented regions or sparse landmarks); 2) there exists a clear gap between multi-view images rendered from textureless 3D shapes and natural images on which LVMs are trained, the feature extracted from LVMs can be noisy; 3) many approaches leverage the semantics in a simple feed-forward manner and therefore require complicated filtering schemes to ensure reliability [202].
@@
Figure 3. Our pipeline mainly contains three blocks. (a) Given two extrinsically aligned shapes, we first render them from multi-views, and then use DiffMorpher [342] to generate image interpolation with respect to each views. (b) We then use SC-GS 3D Gaussian Splatting [329] to reconstruct, from which we obtain a set of dense and noisy point clouds (in gold). They are fed into our surface point extraction module to obtain clean intermediate point clouds (in blue). (c) Finally, we estimate a continuous normalizing flow, represented by an MLP, that deforms the source to the target under the guidance of the extracted intermediate point clouds (in blue).
@@
We believe the above issues originate from the fact that LVMs are used to extract features from shapes in a static, independent manner, which can be problematic when the shapes are distinct from each other. In light of this, we propose SRIF, a framework for Semantic shape Registration, which is built with diffusion-based Image morphing and Flow estimation. The key insight of SRIF is to take a dynamic viewpoint of leveraging LVMs to associate 3D shapes. More specifically, given a pair of extrinsically aligned shapes, we first render them from multi-views, then we utilize a diffusion-based image interpolation framework [342] to generate sequences of intermediate images between corresponding views. Then we reconstruct the intermediate point clouds from the interpolated images via a dynamic 3D Gaussian splatting reconstruction framework [130] and our surface point extractor tailored for the reconstructed Gaussians. Note that, with the above procedures, we have not achieved any explicit connection between the input shapes. Nevertheless, the intermediate point clouds carry rich information on how the shapes are associated from the viewpoint of LVMs. Last but not least, inspired by the dynamic nature of the extracted semantics, we formulate the shape registration problem as estimating a flow that deforms source shape towards the target, with the intermediate point clouds as guidance. In particular, we adopt the framework of PointFlow [324] into our pipeline and learn a Multi Layer Perception (MLP) to represent as well as optimize for the flow.
@@
We extensively evaluate SRIF on a wide range of shape pairs from SHREC’07 dataset [103] and EBCM [328]. Empirical evidence demonstrates that SRIF outperforms the competing baselines in all test sets. As shown in Fig. 1 and Sec. 4, SRIF not only delivers high-quality dense correspondences between shapes but also generates a continuous, semantically meaningful morphing process, which can potentially contribute to 3D data accumulation.
@@ 2.1. Dense Shape Correspondence Estimation
Since estimating 3D shape correspondence is an extensively studied area, we refer readers to [289] for a comprehensive survey and focus on the most relevant methods, which are autonomous methods for general shape matching or registration.
Axiomatic methods typically follow a coarse-to-fine manner, which depends on sparse landmark correspondences to achieve dense ones. The typical approaches [144; 82] first extract and match landmarks using geometric features, and then independently or jointly optimize for dense correspondences based on certain distortion quantity, such as conformality, smoothness. On the other hand, there exists a line of works leveraging fuzzy correspondences [212] to alleviate under-constrained space of dense maps. For instance, MapTree [242] exploits the space of functional maps, SmoothShells [83] jointly estimate registration transform in both spatial and functional space. As mentioned before, correspondences derived from geometry do not always align with semantics, especially in the presence of heterogeneity. Our method enjoys the semantic information from LVMs for accurate correspondence estimation.
@@
1) Large Vision Models (LVMs): SATR [3] renders 3D shapes into multi-view images, gathers coarse semantic labels as part segmentation from each view, and finally back-projects to 3D shapes. Similarly to the axiomatic methods, such cues can be further post-processed with off-the-shelf methods like functional maps [212] to achieve dense correspondences [2]. Since the linguistic signal is not suitable for point-level description, the above approaches suffer from coarse external semantics. It is then highly non-trivial to perform fine-grained shape analysis tasks, including registration. NSSM [202] leverages features from Dinov2 [210] to identify sparse landmark correspondences. Being a two-stage method, NSSM can be fragile regarding the mismatched landmarks. On the other hand, as shown in Fig. 4, SRIF is robust with poor semantic information from LVMs.
2) Category-specific Training is prevailing in learning-based 3D shape analysis. We identify the most relevant approach along this line as NeuroMorph [84], which jointly learns to interpolate and perform registration between shapes. However, NeuroMorph requires pre-training on a small set of 3D shapes, which belongs to the same category as the test ones. Moreover, the interpolation produced by NeuroMorph is dominated by geometric cues, rather than semantics. NeuroMorph is related to our method in the sense that it delivers interpolation as a by-product as well. However, our method generates plausible interpolations before estimating correspondences, while NeuroMorph jointly optimizes for both. As suggested by the experimental results, our scheme makes full use of LVMs and surpasses NeuroMorph by a large margin.
@@ 2.2. Diffusion Model
Diffusion models [117; 254] have gained significant popularity recently, thanks to their impressive capacity of learning data distribution from large-scale image datasets. Some recent works [343; 254; 29] have attempted to control the generated results and improve the quality of generation. On the other hand, the application of diffusion models in image interpolation receives relatively less attention. During the interpolation process using 2D diffusion models, there is a greater focus on the style of the images rather than on deformation. Want et al. [306] attempt to interpolate in the latent space of diffusion, but the resulting method suffers from poor generalization capability. On the other hand, DiffMorpher [342] has achieved smooth interpolation based on StableDiffusion. We therefore exploit it in our pipeline as a tool for image morphing, which returns plausible intermediate images encoding the deformations between the source and the target shape.
@@
As a scene representation, 3D Gaussian Splatting (3D-GS) [138] represents a 3D scene as a mixture of Gaussian distribution. 3D-GS has dramatically advanced novel view synthesis (NVS) in terms of accuracy and efficiency. Numerous follow-ups have been proposed on 3D-GS, ranging across dynamic NVS [329; 130], SLAM [136], and geometry recovery [108], to name a few.
In particular, our method directly leverages SC-GS [130] to reconstruct intermediate geometry from the morphed images. Our point cloud extraction (Sec. 3.2) is similar to mesh extraction from 3D-GS [108], which remains a challenging open problem.
@@ 3.1. Image Rendering and Morphing
The key sub-goal of our method is to infer an intermediate morphing process between input shapes. Our first step is to employ a diffusion-based image morphing technique, DiffMorpher [342] on multi-view image sets, rendered to both $\mathcal{S}$ and $\mathcal{T}$.
Specifically, we pre-process the input shapes such that they are centered around the origin and scaled to be inside a unit sphere. For $\mathcal{S}$ (resp. $\mathcal{T}$), we render $K$ views, where the viewpoints are sampled uniformly around the shape in the sphere space with a radius of 3.5 length units. We use the renderer from Open3D library [346]. We observe that properly endowing texture to shapes plays a critical role in the follow-up image morphing step. That is because diffusion models [254] are generally trained on realistic images, which are distinctive from straightforward renderings of textureless shapes. To this end, we explore various coloring schemes for rendering shapes and settle down at the following, which integrates spatial coordinate information (see more details in the Supp. Mat.). Given a shape, we denote by $z_{\mbox{max}},z_{\mbox{min}}$ respectively the maximal and minimal value of the $z-$coordinates of its points, and formulate function $[C_{R},C_{G},C_{B}]=[\lfloor\frac{d}{256}\rfloor,\lfloor\frac{d}{256}\rfloor,d-256\times\lfloor\frac{d}{256}\rfloor]$ to assign color to each point $(x,y,z)$, where $d=\lfloor\frac{z-z_{\mbox{min}}}{z_{\mbox{max}}-z_{\mbox{min}}}\times 65535\rfloor$ and $C_{R},C_{G},C_{B}$ are respective the intensity of the $RGB$ channels. The rendered images can be represented as a set of image pairs: $\mathcal{C}_{r}=\{(I^{\mathcal{S}}_{i},I^{\mathcal{T}}_{i})|i=1,2,...,K\}$. Subsequently, $\mathcal{C}_{r}$ is processed through an image morphing algorithm DiffMorpher [342], which leverages diffusion models to generate $K$ sequences of intermediate images. This technique allows for a more nuanced and continuous transformation between corresponding views within each image pair in $\mathcal{C}_{r}$. For each pair, DiffMorpher generates $T$ intermediate images, transitioning from $I^{\mathcal{S}}_{i}$ to $I^{\mathcal{T}}_{i}$. Consequently, this process yields $\mathcal{C}=\{I^{j}_{i}\mid i=1,2,\ldots,K;j=1,2,\ldots,T\}$, a comprehensive image set that captures a wide array of views and detailed morphing stages.
@@ 3.2. Intermediate Point Clouds Reconstruction and Post-processing
Therefore, we instead formulate the reconstruction as a dynamic one. Moreover, for the sake of efficiency and accuracy, we choose the recent art, SC-GS framework [130], for reconstruction. Specifically, we use the vertices of the source mesh to initialize the spatial position of each Gaussian (i.e., mean). From these points, SC-GS creates a set of 3D Gaussians $G(x,r,s,\sigma)$ defined by a center position $x$, opacity $\sigma$, and 3D covariance matrix $\Sigma$ obtained from quaternion $r$ and scaling $s$. For a morphing stage $t\in\{1,2,\cdots,T\}$, SC-GS takes the positions $x$ as input and predict the $(\delta x,\delta r,\delta s)$. Subsequently, the deformed 3D Gaussians $G(x+\delta x,r+\delta r,s+\delta s,\sigma)$ is optimized by the differential Gaussian rasterization pipeline. Once optimization is done, given a time step $t$, we extract the set of positions as the raw intermediate point clouds $V_{t}^{G}$.
@@
Though it seems natural to iteratively perform shape registration [12] between consecutive point clouds in $\mathcal{V}$, we observe that this naive approach can lead to sub-optimal results. As shown in the top row of Fig. 4, due to the significant deformation between the cow and the giraffe, the intermediate point clouds are of low quality. Deforming the cow towards such can lead to distortion accumulation (see the middle row of Fig. 4).
@@
In addition, to ensure the shape does not undergo excessive distortion during the registration process, we include an ARAP (As-Rigid-As-Possible) loss followed by [111] as a regularization term. The total cost function $E_{\text{total }}$ combines the above terms with the weighting factors $\lambda_{\mbox{cd\_inter}},\lambda_{\mbox{cd\_final}}$, and $\lambda_{\mbox{arap}}$ to balance them:
@@
To achieve more accurate registration results, we perform non-rigid shape registration [12] between the output of the flow network and $\mathcal{T}$.

### v2

@@ Abstract.
ccs: Computing methodologies Shape analysis
@@ 1. Introduction
Estimating dense correspondences between 3D shapes serves as a cornerstone in many applications of computer graphics, including 3D reconstruction (Yu et al., 2018), animation (Sumner and Popović, 2004) and statistical shape analysis (Anguelov et al., 2005), to name a few. Regarding shapes undergoing rigid or isometric deformations, the prior shape registration/matching techniques (PJ and ND, 1992; Amberg et al., 2007; Bronstein et al., 2006; Ovsjanikov et al., 2012) have laid down solid foundations on both theoretical and practical fronts. In this paper, we consider the problem of estimating semantically meaningful dense correspondences between shapes undergoing more general and complicated deformations.
@@
In the absence of a compact deformation prior, the purely geometric methods typically take a coarse-to-fine approach. Namely, one leverages geometric features to locate a small set of landmarks on both shapes, estimates sparse landmark correspondences, and finally propagates dense correspondences via minimizing distortions such as conformal (Kim et al., 2011), elastic energy (Edelstein et al., 2020). It is worth noting, though, that the sparse correspondences derived from geometry are not necessarily relevant to semantics. As shown in qualitative results in Sec. 4, the resulting maps can suffer from such discrepancy, especially in the presence of significant heterogeneity.
@@
To this end, another line of works concentrate on producing high-quality semantic correspondences at the cost of dependency on user-defined landmarks (Aigerman et al., 2015; Mandad et al., 2017; Ezuz et al., 2019; Schmidt et al., 2023; Yang et al., 2020). Treating landmarks as anchor points, they cast the problem as correspondence interpolation, which can be conducted by optimizing certain geometric distortions. However, the dependency on manual annotations limits the practical utility of these works. Apart from hindering automation, insufficient annotation strength can lead to sub-optimal results as shown in Fig. 2.
@@
Figure 2. Maps obtained by state-of-the-art compatible remeshing techniques (Yang et al., 2020; Schmidt et al., 2023) based on varying number of landmark correspondences. The absence of landmarks at local region can lead to erroneous matching (see the red boxes). On the other hand, our method delivers high-quality maps with no landmark annotation.
@@
More recently, the emergence of learning-based techniques has enabled data-driven semantic information extraction. In essence, one can learn canonical structural information from a collection of 3D shapes, which can help to match challenging non-rigid shapes (Sun et al., 2021; Eisenberger et al., 2021), or heterogeneous man-made objects (Deng et al., 2021; Zheng et al., 2021). Unfortunately, limited by the amount of available 3D data, the above approaches are typically category-specific, weakening their practical utility.
@@
On the other hand, boosted by large-scale natural image datasets such as LAION (Schuhmann et al., 2022), large vision (or vision-language) models (LVMs)  (Oquab et al., 2023; Rombach et al., 2022; Radford et al., 2021; OpenAI, 2023) has attracted considerable attention from the community of 3D shape analysis  (Abdelreheem et al., 2023b, a; Morreale et al., 2024; Wimmer et al., 2024). The shared protocol is to project 3D shapes into multi-view 2D images. The latter can then be fed into LVMs for semantic encoding, which is finally aggregated back to 3D shapes for the respective task. While the above approach has been proven simple yet effective, we observe that 1) the distilled semantic information is in general coarse (e.g., segmented regions or sparse landmarks); 2) there exists a clear gap between multi-view images rendered from textureless 3D shapes and natural images on which LVMs are trained, the feature extracted from LVMs can be noisy; 3) many approaches leverage the semantics in a simple feed-forward manner and therefore require complicated filtering schemes to ensure reliability (Morreale et al., 2024).
@@
Figure 3. Our pipeline mainly contains three blocks. (a) Given two extrinsically aligned shapes, we first render them from multi-views, and then use DiffMorpher (Zhang et al., 2023) to generate image interpolation with respect to each views. (b) We then use SC-GS 3D Gaussian Splatting (Yang et al., 2023) to reconstruct, from which we obtain a set of dense and noisy point clouds (in gold). They are fed into our surface point extraction module to obtain clean intermediate point clouds (in blue). (c) Finally, we estimate a continuous normalizing flow, represented by an MLP, that deforms the source to the target under the guidance of the extracted intermediate point clouds (in blue).
@@
We believe the above issues originate from the fact that LVMs are used to extract features from shapes in a static, independent manner, which can be problematic when the shapes are distinct from each other. In light of this, we propose SRIF, a framework for Semantic shape Registration, which is built with diffusion-based Image morphing and Flow estimation. The key insight of SRIF is to take a dynamic viewpoint of leveraging LVMs to associate 3D shapes. More specifically, given a pair of extrinsically aligned shapes, we first render them from multi-views, then we utilize a diffusion-based image interpolation framework (Zhang et al., 2023) to generate sequences of intermediate images between corresponding views. Then we reconstruct the intermediate point clouds from the interpolated images via a dynamic 3D Gaussian splatting reconstruction framework (Huang et al., 2024) and our surface point extractor tailored for the reconstructed Gaussians. Note that, with the above procedures, we have not achieved any explicit connection between the input shapes. Nevertheless, the intermediate point clouds carry rich information on how the shapes are associated from the viewpoint of LVMs. Last but not least, inspired by the dynamic nature of the extracted semantics, we formulate the shape registration problem as estimating a flow that deforms source shape towards the target, with the intermediate point clouds as guidance. In particular, we adopt the framework of PointFlow (Yang et al., 2019) into our pipeline and learn a Multi Layer Perception (MLP) to represent as well as optimize for the flow.
@@
We extensively evaluate SRIF on a wide range of shape pairs from SHREC’07 dataset (Giorgi et al., 2007) and EBCM (Yang et al., 2020). Empirical evidence demonstrates that SRIF outperforms the competing baselines in all test sets. As shown in Fig. 1 and Sec. 4, SRIF not only delivers high-quality dense correspondences between shapes but also generates a continuous, semantically meaningful morphing process, which can potentially contribute to 3D data accumulation.
@@ 2.1. Dense Shape Correspondence Estimation
Since estimating 3D shape correspondence is an extensively studied area, we refer readers to (Tam et al., 2013) for a comprehensive survey and focus on the most relevant methods, which are autonomous methods for general shape matching or registration.
Axiomatic methods typically follow a coarse-to-fine manner, which depends on sparse landmark correspondences to achieve dense ones. The typical approaches (Kim et al., 2011; Edelstein et al., 2020) first extract and match landmarks using geometric features, and then independently or jointly optimize for dense correspondences based on certain distortion quantity, such as conformality, smoothness. On the other hand, there exists a line of works leveraging fuzzy correspondences (Ovsjanikov et al., 2012) to alleviate under-constrained space of dense maps. For instance, MapTree (Ren et al., 2020) exploits the space of functional maps, SmoothShells (Eisenberger et al., 2020) jointly estimate registration transform in both spatial and functional space. As mentioned before, correspondences derived from geometry do not always align with semantics, especially in the presence of heterogeneity. Our method enjoys the semantic information from LVMs for accurate correspondence estimation.
@@
1) Large Vision Models (LVMs): SATR (Abdelreheem et al., 2023b) renders 3D shapes into multi-view images, gathers coarse semantic labels as part segmentation from each view, and finally back-projects to 3D shapes. Similarly to the axiomatic methods, such cues can be further post-processed with off-the-shelf methods like functional maps (Ovsjanikov et al., 2012) to achieve dense correspondences (Abdelreheem et al., 2023a). Since the linguistic signal is not suitable for point-level description, the above approaches suffer from coarse external semantics. It is then highly non-trivial to perform fine-grained shape analysis tasks, including registration. NSSM (Morreale et al., 2024) leverages features from Dinov2 (Oquab et al., 2023) to identify sparse landmark correspondences. Being a two-stage method, NSSM can be fragile regarding the mismatched landmarks. On the other hand, as shown in Fig. 4, SRIF is robust with poor semantic information from LVMs.
2) Category-specific Training is prevailing in learning-based 3D shape analysis. We identify the most relevant approach along this line as NeuroMorph (Eisenberger et al., 2021), which jointly learns to interpolate and perform registration between shapes. However, NeuroMorph requires pre-training on a small set of 3D shapes, which belongs to the same category as the test ones. Moreover, the interpolation produced by NeuroMorph is dominated by geometric cues, rather than semantics. NeuroMorph is related to our method in the sense that it delivers interpolation as a by-product as well. However, our method generates plausible interpolations before estimating correspondences, while NeuroMorph jointly optimizes for both. As suggested by the experimental results, our scheme makes full use of LVMs and surpasses NeuroMorph by a large margin.
@@ 2.2. Diffusion Model
Diffusion models (Ho et al., 2020; Rombach et al., 2022) have gained significant popularity recently, thanks to their impressive capacity of learning data distribution from large-scale image datasets. Some recent works (Zhao et al., 2023; Rombach et al., 2022; Blattmann et al., 2023) have attempted to control the generated results and improve the quality of generation. On the other hand, the application of diffusion models in image interpolation receives relatively less attention. During the interpolation process using 2D diffusion models, there is a greater focus on the style of the images rather than on deformation. Want et al. (Wang and Golland, 2023) attempt to interpolate in the latent space of diffusion, but the resulting method suffers from poor generalization capability. On the other hand, DiffMorpher (Zhang et al., 2023) has achieved smooth interpolation based on StableDiffusion. We therefore exploit it in our pipeline as a tool for image morphing, which returns plausible intermediate images encoding the deformations between the source and the target shape.
@@
As a scene representation, 3D Gaussian Splatting (3D-GS) (Kerbl et al., 2023) represents a 3D scene as a mixture of Gaussian distribution. 3D-GS has dramatically advanced novel view synthesis (NVS) in terms of accuracy and efficiency. Numerous follow-ups have been proposed on 3D-GS, ranging across dynamic NVS (Yang et al., 2023; Huang et al., 2024), SLAM (Keetha et al., 2023), and geometry recovery (Guédon and Lepetit, 2023), to name a few.
In particular, our method directly leverages SC-GS (Huang et al., 2024) to reconstruct intermediate geometry from the morphed images. Our point cloud extraction (Sec. 3.2) is similar to mesh extraction from 3D-GS (Guédon and Lepetit, 2023), which remains a challenging open problem.
@@ 3.1. Image Rendering and Morphing
The key sub-goal of our method is to infer an intermediate morphing process between input shapes. Our first step is to employ a diffusion-based image morphing technique, DiffMorpher (Zhang et al., 2023) on multi-view image sets, rendered to both $\mathcal{S}$ and $\mathcal{T}$.
Specifically, we pre-process the input shapes such that they are centered around the origin and scaled to be inside a unit sphere. For $\mathcal{S}$ (resp. $\mathcal{T}$), we render $K$ views, where the viewpoints are sampled uniformly around the shape in the sphere space with a radius of 3.5 length units. We use the renderer from Open3D library (Zhou et al., 2018). We observe that properly endowing texture to shapes plays a critical role in the follow-up image morphing step. That is because diffusion models (Rombach et al., 2022) are generally trained on realistic images, which are distinctive from straightforward renderings of textureless shapes. To this end, we explore various coloring schemes for rendering shapes and settle down at the following, which integrates spatial coordinate information (see more details in the Supp. Mat.). Given a shape, we denote by $z_{\mbox{max}},z_{\mbox{min}}$ respectively the maximal and minimal value of the $z-$coordinates of its points, and formulate function $[C_{R},C_{G},C_{B}]=[\lfloor\frac{d}{256}\rfloor,\lfloor\frac{d}{256}\rfloor,d-256\times\lfloor\frac{d}{256}\rfloor]$ to assign color to each point $(x,y,z)$, where $d=\lfloor\frac{z-z_{\mbox{min}}}{z_{\mbox{max}}-z_{\mbox{min}}}\times 65535\rfloor$ and $C_{R},C_{G},C_{B}$ are respective the intensity of the $RGB$ channels. The rendered images can be represented as a set of image pairs: $\mathcal{C}_{r}=\{(I^{\mathcal{S}}_{i},I^{\mathcal{T}}_{i})|i=1,2,...,K\}$. Subsequently, $\mathcal{C}_{r}$ is processed through an image morphing algorithm DiffMorpher (Zhang et al., 2023), which leverages diffusion models to generate $K$ sequences of intermediate images. This technique allows for a more nuanced and continuous transformation between corresponding views within each image pair in $\mathcal{C}_{r}$. For each pair, DiffMorpher generates $T$ intermediate images, transitioning from $I^{\mathcal{S}}_{i}$ to $I^{\mathcal{T}}_{i}$. Consequently, this process yields $\mathcal{C}=\{I^{j}_{i}\mid i=1,2,\ldots,K;j=1,2,\ldots,T\}$, a comprehensive image set that captures a wide array of views and detailed morphing stages.
@@ 3.2. Intermediate Point Clouds Reconstruction and Post-processing
Therefore, we instead formulate the reconstruction as a dynamic one. Moreover, for the sake of efficiency and accuracy, we choose the recent art, SC-GS framework (Huang et al., 2024), for reconstruction. Specifically, we use the vertices of the source mesh to initialize the spatial position of each Gaussian (i.e., mean). From these points, SC-GS creates a set of 3D Gaussians $G(x,r,s,\sigma)$ defined by a center position $x$, opacity $\sigma$, and 3D covariance matrix $\Sigma$ obtained from quaternion $r$ and scaling $s$. For a morphing stage $t\in\{1,2,\cdots,T\}$, SC-GS takes the positions $x$ as input and predict the $(\delta x,\delta r,\delta s)$. Subsequently, the deformed 3D Gaussians $G(x+\delta x,r+\delta r,s+\delta s,\sigma)$ is optimized by the differential Gaussian rasterization pipeline. Once optimization is done, given a time step $t$, we extract the set of positions as the raw intermediate point clouds $V_{t}^{G}$.
@@
Though it seems natural to iteratively perform shape registration (Amberg et al., 2007) between consecutive point clouds in $\mathcal{V}$, we observe that this naive approach can lead to sub-optimal results. As shown in the top row of Fig. 4, due to the significant deformation between the cow and the giraffe, the intermediate point clouds are of low quality. Deforming the cow towards such can lead to distortion accumulation (see the middle row of Fig. 4).
@@
In addition, to ensure the shape does not undergo excessive distortion during the registration process, we include an ARAP (As-Rigid-As-Possible) loss followed by (Guo et al., 2021) as a regularization term. The total cost function $E_{\text{total }}$ combines the above terms with the weighting factors $\lambda_{\mbox{cd\_inter}},\lambda_{\mbox{cd\_final}}$, and $\lambda_{\mbox{arap}}$ to balance them:
@@
To achieve more accurate registration results, we perform non-rigid shape registration (Amberg et al., 2007) between the output of the flow network and $\mathcal{T}$.